OpenStudy (anonymous):

Find the radius of a circle whose area equals the area of a rectangle that measures 2 ft. by 11 ft. (Use = 22/7)

OpenStudy (anonymous):

area of circle = area of rectangle \[\pi \times r ^{2}=length \times breadth\] (22/7) * r^2 = 2 * 11